### v3

Abstract.
Efficient scene representations are essential for many computer graphics applications. A general unified representation that can handle both surfaces and volumes simultaneously, remains a research challenge. In this work we propose a compact and efficient alternative to existing volumetric representations for rendering such as voxel grids. Inspired by recent methods for scene reconstruction that leverage mixtures of 3D Gaussians to model radiance fields, we formalize and generalize the modeling of scattering and emissive media using mixtures of simple kernel-based volumetric primitives. We introduce closed-form solutions for transmittance and free-flight distance sampling for different kernels, and propose several optimizations to use our method efficiently within any off-the-shelf volumetric path tracer. We demonstrate our method in both forward and inverse rendering of complex scattering media. Furthermore, we adapt and showcase our method in radiance field optimization and rendering, providing additional flexibility compared to current state of the art given its ray-tracing formulation. We also introduce the Epanechnikov kernel and demonstrate its potential as an efficient alternative to the traditionally-used Gaussian kernel in scene reconstruction tasks. The versatility and physically-based nature of our approach allows us to go beyond radiance fields and bring to kernel-based modeling and rendering any path-tracing enabled functionality such as scattering, relighting and complex camera models. Code is available at https://github.com/facebookresearch/volumetric_primitives
Volume Rendering, Scattering, Radiance Fields, 3D Reconstruction, Volumetric Primitives, Volumetric Representations, Ray Tracing, Inverse Rendering
copyright: ccjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 1article: 10publicationmonth: 2doi: 10.1145/3711853ccs: Computing methodologies Shape representationsccs: Computing methodologies Ray tracingccs: Computing methodologies Volumetric modelsccs: Computing methodologies Point-based modelsccs: Computing methodologies Reconstruction
1. Introduction
Volumetric representations of appearance have found good success at representing complex appearances such as cloth and hair (shroder11; zhao11; khungurn2015; aliaga17), trees (neyret98; loubet2017), clouds and smoke (kallweit2017), or particle aggregates (moon07; meng15; mueller16efficient), to name a few. In this work we propose a compact volumetric representation of appearance, that allows efficient physics-based rendering of both scattering and emissive media.
Most previous approaches for modeling volumetric appearances leverage grid-like representations of appearance such as voxels-grids, potentially with a multilevel underlying hierarchy, for which highly efficiently libraries exist (museth2013vdb). Voxel grids are standard in rendering, and used in most research (jakob2010mitsuba; jakob2022mitsuba3; pbrt) and production renderers (fajardo18arnold; renderman; hyperion; manuka). Voxel-grids are flexible, provide $O(1)$ query time, and spatial connectivity for e.g., downsampling and level-of-detail representations (heitz2015sggx; loubet2017). Unfortunately, they scale poorly in terms of memory, specially when representing very fine detail and sparse structures. While hierarchical grid-based methods (ProductionVolumeRendering2017) have proven very effective at subdividing complex heterogeneous models into simpler local components, their building blocks are still grid-based, essentially inheriting the same problems. For an efficient integration of e.g., transmittance, voxel grids require the use of stochastic tracking techniques (novak2018monte), which introduces additional variance when rendering media.
In parallel, volume rendering techniques and continuous volumetric representations have recently seen unprecedented interest in the fields of computer vision and image-based graphics, spearheaded by efforts such as neural radiance fields (mildenhall2020). These methods allow to capture and render photorealistic three-dimensional scenes, by optimizing an underlying volumetric representation of matter. Different volumetric representations have been proposed, including point-based methods (kerbl20233Dgaussians; sainz04), implicit neural models (mildenhall2020; mipnerf; barron2022mipnerf360), explicit voxel-based representations (yu21plenoxels; yu2021plenoctrees), and hybrid approaches (mueller2022instant; pointnerf; Lombardi21). While implicit neural models generally achieve the highest compression rates, point-based methods can be considered as the state of the art in terms of image quality and rendering speed for radiance fields rendering, while still being substantially more compact than voxel-grid structures. However, while their derivation starts from physics (i.e., the volume rendering equation), these methods in general are not suitable for their use within a physics-based renderer, given the fairly large assumptions and simplifications imposed in their image formation model to favour speed and ease of optimization.
Inspired by the success of recent 3D Gaussian-based representations of radiance fields (kerbl20233Dgaussians), we propose a new representation for general scattering and emissive media based on mixtures of three-dimensional kernel-based volumetric primitives. Each primitive statistically represents a spatial distribution of matter with the same optical properties. We introduce this representation inside the radiative transfer theory (RTT) (chandrasekhar1960radiative) and derive closed-form solutions for transmittance and emission without the need of stochastic sampling, as well as sampling routines for computing inscattering.
While the scene representation is similar to the Gaussian splatting technique (kerbl20233Dgaussians), the image formation model is fundamentally different: Kerbl et al. assume that each primitive becomes a billboard oriented towards the camera and therefore the density along the ray becomes a sum of delta functions; this allows them to implement their model in an extremely efficient rasterization pipeline. In contrast, our radiative primitive-based formulation preserves the primitive’s three-dimensional density and integrates following the physical transport process. This is crucial for physics-based rendering since it allows preserving reciprocity, while solving some view-dependent primitive ordering problems in Gaussian splatting. Our theoretical framework is general and supports a wide range of three-dimensional kernels. We demonstrate our work using the Gaussian kernel employed by Kerbl et al. (kerbl20233Dgaussians), but also on 3D Epanechnikov kernels, implementing closed-form solutions for volumetric transmittance, emission, and free-flight sampling. We implement our work by leveraging hardware-accelerated ray tracing for finding the primitives’ bounds and integrating their contribution along the ray. This fits well inside modern ray-tracing-based renderers, allowing us to compute transmittance and emission for both primary and secondary rays, needed for multiple scattering. We also derive the adjoint of our image formation model, which we demonstrate in inverse rendering applications. We showcase our primitive-based volumetric model through several applications both in forward and inverse rendering: 1) traditional forward rendering of scattering media, where we demonstrate improved performance over voxel-grids using local aggregated statistics and state of the art transmittance estimators; 2) inverse rendering of scattering and purely absorptive media; 3) inverse tomographic reconstructions from focus stacks using telecentric cameras and 4) radiance field optimization and rendering of complex scenes (both real and synthetic), with increased generality and controllability. Furthermore, our integration into the general RTT allows for future extensions in relighting and scattering. In short, our contributions are:
  • a novel kernel-based primitive representation for volumetric media that fits into the radiative transfer framework and can be integrated in any physics-based rendering engine;
  • closed-form expressions for transmittance and emission, as well as distance sampling routines proportional to transmittance, for 3D Gaussian and Epanechnikov kernels;
  • an efficient ray tracing-based implementation for solving light transport using our novel volumetric representation for both scattering and emissive (radiance field) media;
  • and adjoint derivatives of our forward methods for solving inverse reconstruction problems efficiently.
2. Related Work
Volumetric light transport
Simulating light transport in media has been thoroughly studied in computer graphics (novak2018monte). It involves solving the radiative transfer equation (RTE) (chandrasekhar1960radiative), which has been extended to account for anisotropic (jakob2010radiative), refractive (ament14refractive), or spatially correlated (jarabo2018radiative; bitterli2018radiative) media. These generalizations make it possible to render a wider range of light transport phenomena.
Numerical estimation of transmittance is at the core of modern volumetric light transport simulation (novak2018monte); via biased quadrature rules (ray-marching) (tuy1984direct; munoz2014higher) and modern unbiased variants (kettunen21), to stochastic tracking algorithms (null-scattering estimators) (woodcock), including extensions with variance reduction via control variates (novak14; szirmay2017unbiased; crespo2021primary), alternative formulations based on power-series (georgiev19), or differentiable formulations targeting inverse rendering (nimierdavid2022unbiased). All these approaches require tight estimates of the maximum density (majorant) for performance, either obtained in precomputation (yue2010unbiased; szirmay2011free) or estimated on-the-fly (carter1972monte; kutz2017spectral; galtier2013integral; misso2023progressive). Our approach, on the other hand, allows to compute transmittance in closed-form, as the product of transmittance of all primitives along the ray.
Volumetric representations of matter
For heterogeneous media, discrete voxel-grid approaches are ubiquitously found in most applications due to their simplicity and flexibility. They can feature multi-level hierarchical structures (museth2013vdb; museth2021nanovdb), which are helpful for faster traversal and filtering for level-of-detail applications, but suffer from poor scalability in terms of memory consumption when the modeled medium possesses finer details. Sparser representations like collections of isotropic and simple volumes have been used in other fields, e.g., particle physics (Brown03), where they can model neutron transport in graphite pebble-bed reactors. The advantage of having collections of these simple volumes is that, individually, they offer closed-form solutions for transmittance estimation and sampling, reducing variance when compared with most tracking methods normally used in voxel grids. However, searching for the boundaries of these volumes can be inefficient and slow in many situations (bitterli2018radiative). In graphics, we can find hierarchical grid-based approaches (ProductionVolumeRendering2017) for subdividing complex, potentially sparse, volumes into smaller, more compact representations, though eventually they rely on smaller voxel-based primitives. Alternatively, implicit representations of participating media have been proposed, most notably in the context of radiance fields but also in compression of large volumes, including large multilayer perceptrons (MLP) (mildenhall2020; neuralvdb) or sparse hash grids of features combined with tiny MLPs (mueller2022instant). These approaches result in large compression rates, at the cost of expensive queries. Closer to our work, Knoll et al. (Knoll2021) proposed to model emissive and absorbing media using isotropic Gaussian mixture models, in the context of particle-based volumes (max1979atomlll); however, they did not leverage closed-forms expressions for transmittance, requiring stochastic integration via tracking, and cannot represent scattering media. In contrast, our work proposes closed-form solutions for transmittance and sampling, and further extends to support scattering media, anisotropic kernels, and efficient inverse rendering.
Inverse volumetric rendering
Early works on inverse rendering of heterogeneous media used inverse volumetric rendering for matching micron-scale cloth patches to measures (khungurn2015), approximating multiscale volumetric representations (zhao2016downsample), scattering compensation for 3D printing (elek17; sumin19; nindel2021gradient; Condor2023), or inverse scattering (Gkioulekas2016AFF) by using hand-made derivatives of light transport. Zhang et al. (Zhang2019drt) proposed a differential radiative transfer framework suitable for inverse volumetric rendering. These works use voxel grids for representing media density, though our proposed representation would fit in these inverse pipelines.
Radiance fields
Neural Radiance Fields (NeRF) (mildenhall2020) introduced a continuous, implicit volumetric representation based on multi-layer perceptrons (MLPs). Subsequent works have leveraged multi-scale representations (mipnerf) and extended to unbounded scenes (barron2022mipnerf360; zhang2020nerf++; reiser2023merf); however, training and evaluating the MLPs still requires significant computation. In contrast, Instant-NGP (mueller2022instant) relies on a sparse hash grid of features and a small MLP, enabling faster scene reconstruction. Sparsification and hierarchization has been a popular avenue for optimizing radiance field fitting and rendering (yu2021plenoctrees; hedman2021snerg; bovzivc2022neu=ral; chen2022mobilenerf; Lombardi21), achieving real-time rendering even on mobile devices, or helping on accelerating off-line rendering with complex assets (zhu2021neural; condor22). Most recently, 3D Gaussian Splatting (kerbl20233Dgaussians) removes the need for MLPs altogether, relying instead on an explicit Gaussian-based representation. Combined with an efficient splatting technique for rasterization (ewasplattingzwicker04), their approach is able to achieve fast rendering performance and high-quality reconstruction, allocating more primitives in areas where finer detail is needed. Our work takes Kerbl’s approach, and extends it to general scattering media in the context of the radiative transfer framework.
Gaussians in physically-based rendering
Gaussian mixture models defined over the sphere have been extensively used in graphics for fitting BRDFs or environment maps (Xu13sigasia), precomputed radiance transfer (green2006view), or on-line path guiding (vorba2014line). Yan et al. (yan2016position) fitted a 4D Gaussian mixture model for accelerating computation when rendering glinty surfaces. Jakob et al. (jakob2011progressive) proposed to fit a 3D anisotropic Gaussian mixture model to the distribution of radiance in media, showing accelerated rendering using photon-based techniques. Despite their work sharing some similarities to ours (closed-form integration along Gaussians), it focuses on fitting radiance in the context of radiance estimation, while our approach focuses on modeling matter (density) and can be used in arbitrary volumetric renderers. Combining both representations (kernel-based representations for both media and radiance) would be an interesting line of future work.
3. Background
Under the geometric optics assumption, light transport in participating media is governed by the radiative transfer equation (RTE) (chandrasekhar1960radiative). We compute the incident radiance $L(\mathbf{x},\omega)$ at point $\mathbf{x}$ and direction $\omega$ as
(1) $(\omega\cdot\nabla)L(\mathbf{x},\omega)$ $=-\mu_{a}(\mathbf{x})L(\mathbf{x},\omega)-\mu_{s}(\mathbf{x})L(\mathbf{x},\omega)$
$+\mu_{a}(\mathbf{x})Q(\mathbf{x},\omega)+\mu_{s}(\mathbf{x})S(\mathbf{x},\omega),$
where $S$ is the in-scattering term, defined as
(2) $S(\mathbf{x},\omega)=\int_{\mathcal{S}^{2}}f_{\text{p}}(\mathbf{x},\omega^{\prime}\rightarrow\omega)L(\mathbf{x},\omega^{\prime})\,\text{d}\omega^{\prime},$
that models the light scattered in direction $\omega$ as a function of the integral of the light incoming from all directions on the unit sphere $\mathcal{S}^{2}$}. The interaction of light and matter is characterized by the optical properties of the medium, described in terms of the differential probabilities of absorption and scattering (or coefficients) $\mu_{a}(\mathbf{x})$ and $\mu_{s}(\mathbf{x})$ respectively, the extinction coefficient $\mu_{t}(\mathbf{x})=\mu_{a}(\mathbf{x})+\mu_{s}(\mathbf{x})$, the phase function $f_{\text{p}}(\mathbf{x},\omega^{\prime}\rightarrow\omega)$, and the emission $Q(\mathbf{x},\omega)$. Assuming uncorrelated statistics of matter, the differential probability of extinction is defined as the product between the density $\rho(\mathbf{x})$ of the particles forming the medium at point $\mathbf{x}$ [m-3] and the particles’ cross section $\sigma(\mathbf{x})$ [m2] as $\mu_{t}(\mathbf{x})=\rho(\mathbf{x})\,\sigma(\mathbf{x})$. For simplicity, we ignore the wavelength dependency in Equation (1), assume scalar elastic light-matter interactions, and omit the potential anisotropy of the medium (jakob2010radiative); generalizing the following derivations to include these additional dimensions is trivial. By integrating both sides of the differential RTE (1) along the direction $\omega$ we obtain the volumetric rendering equation, that models radiance $L(\mathbf{x},\omega)$ as
$L(\mathbf{x}_{0},\omega)$ $=\int_{0}^{s}\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})\,\mu_{t}(\mathbf{x}_{t})\,L_{o}(\mathbf{x}_{t},\omega)\,\text{d}t,$
(3) $+\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{s})L(\mathbf{x}_{s},\omega)\qquad\text{ with }$
(4) $L_{o}(\mathbf{x},\omega)$ $=\left(1-\alpha(\mathbf{x})\right)Q(\mathbf{x},\omega)+\alpha(\mathbf{x})S(\mathbf{x},\omega),$
with $\mathbf{x}_{t}=\mathbf{x}_{0}-t\omega$ and distance $t$, $\alpha(\mathbf{x})=\mu_{s}(\mathbf{x})/\mu_{t}(\mathbf{x})$ as the single scattering albedo, and $\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$ is the transmittance defined as the fractional visibility due to absorption and out-scattering in the medium, and modeled following the Beer-Lambert law as
(5) $\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})=\exp\left(-\int_{0}^{t}\mu_{t}(\mathbf{x}_{t^{\prime}})\text{d}t^{\prime}\right).$
Finally, $L(\mathbf{x}_{s},\omega)$ is the outgoing radiance at the medium boundary $\mathbf{x}_{s}$ at distance $s$ modeled by the rendering equation (kajiya86).
4. Volumetric Primitives
Inspired by recent works using Gaussian primitives for reconstructing radiance fields (kerbl20233Dgaussians), and following a similar medium definition as Knoll et al. (Knoll2021), we model media using sets of primitives. Each primitive $\mathcal{P}_{i}$ represents a statistical aggregate of matter with identical emission, cross section and phase function, and with density $\rho_{i}(\mathbf{x})$ defined by a three-dimensional un-normalized kernel $K_{i}(\mathbf{x})$. Given these primitives, we can model the distribution of matter, and therefore the extinction probability as a mixture of these volumetric primitives following
(6) $\mu_{t}(\mathbf{x})=\sum_{i=1}^{N}\sigma_{i}\,K_{i}(\mathbf{x}),$
with $N$ the number of primitives affecting $\mathbf{x}$, and $\sigma_{i}$ their cross-section. For kernels with infinite support, $N$ is the total number of primitives $N_{\text{total}}=N$. In our case, we assume that all kernels have limited support, so typically $N<N_{\text{total}}$. We compute the single-scattering albedo and phase function at $\mathbf{x}$ analogously, with the key difference of requiring normalization (see Appendix LABEL:app:opt_params).
Figure 2. Top: Flatland representation of a medium modeled using $N_{\text{total}}=5$ primitives $\mathcal{P}_{i}$ with $i\in[1,5]$. For the depicted ray from the camera only three primitives contribute directly to the ray. Bottom: The primitives are projected in 1D along the ray, and their density, emission and in-scattering are integrated over that line, which can be separated in disjoint segments defined by the boundaries of the primitives. The contribution of each segment $\mathcal{S}_{k}$ is the integral over the primitives overlapping on that segment.
[…]
[¡short description¿]¡long description¿
4.1. The RTE with volumetric primitives
Defining matter as a mixture of volumetric primitives with finite support allow us to reinterpret Equation (3) as the summation of the contribution of all the primitives. For that, as shown in Figure 2 (bottom), we split the ray in segments based on the primitives’ entry and exit points; then, we integrate radiance along segments in a front-to-back pass, keeping track of the multiplicative transmittance which is computed analytically. This approach is similar to how production renderers deal with complex overlapping media (ProductionVolumeRendering2017, Ch.6); the key difference is that our approach approximates the media with primitives, which allows having closed-form transmittance and sampling expressions.
More formally, the boundaries of each primitive $t_{i,0}$ and $t_{i,1}$ subdivide the ray into ordered segments $\mathcal{S}_{k}=[\widehat{t}_{k,0},\widehat{t}_{k,1}]$, with $\widehat{t}_{k,0}=\widehat{t}_{k-1,1}$ for $k>0$, with $\widehat{t}_{0,0}=\mathbf{x}$. Each of these segments $\mathcal{S}_{k}$ might overlap with zero, one, or multiple primitives denoted with the set $\{\mathcal{P}_{i}|i\in S_{k}\}$, where $S_{k}$ is the per-segment set of indices, with $|S_{k}|\leq N$ the number of primitives overlapping in segment $\mathcal{S}_{k}$. With that definition of segments along the ray, we rewrite Equation (3) (omitting the boundary condition at $\mathbf{x}_{s}$ for simplicity, as
(7) $L(\mathbf{x}_{0},\omega)=\sum_{k=1}^{M}\textrm{T}_{k-1}(\mathbf{x}_{0},\mathbf{x}_{t})L_{k}(\mathbf{x}_{\widehat{t}_{k,0}},\omega),$
where $M$ is the number of segments along the ray, $L_{k}(\mathbf{x}_{\widehat{t}_{k,0}},\omega)$ the outgoing radiance at the segment $\mathcal{S}_{k}$ defined as
(8) $L_{k}(\mathbf{x}_{\widehat{t}_{k,0}},\omega)=\int_{\widehat{t}_{k,0}}^{\widehat{t}_{k,1}}\textrm{T}(\mathbf{x}_{\widehat{t}_{k,0}},\mathbf{x}_{t})\,\sum_{i\in S_{k}}\left[\sigma_{i}K_{i}(\mathbf{x}_{t})\,L_{o,i}(\mathbf{x}_{t},\omega)\right]\,\text{d}t,$
with $L_{o,i}(\mathbf{x},\omega)$ the outgoing radiance from primitive $\mathcal{P}_{i}$, and $\textrm{T}_{k-1}(\mathbf{x}_{0},\mathbf{x}_{t})$ is the transmittance defined as a recursive operator
(9) $\textrm{T}_{k}(\mathbf{x}_{0},\mathbf{x}_{t})=\textrm{T}_{k-1}(\mathbf{x}_{0},\mathbf{x}_{\widehat{t}_{k-1,1}})\,\textrm{T}(\mathbf{x}_{\widehat{t}_{k,0}},\mathbf{x}_{t}),$
with $\textrm{T}_{0}(\mathbf{x}_{0},\mathbf{x}_{t})=1$ and $\textrm{T}_{M}(\mathbf{x}_{0},\mathbf{x}_{t})=\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$. Trivially, transmittance in segment $k$ is computed as the product of transmittance of all primitives overlapping the segment
(10) $\textrm{T}(\mathbf{x}_{\widehat{t}_{k,0}},\mathbf{x}_{t})=\exp\left(-\sum_{i\in S_{k}}\tau_{i}\left(\mathbf{x}_{\widehat{t}_{k,0}},\mathbf{x}_{\min(t,\widehat{t}_{k,1})}\right)\right),$
with $\tau_{i}(\mathbf{x}_{a},\mathbf{x}_{b})$ the optical depth from primitive $\mathcal{P}_{i}$ in the range $t\in[a,b]$ defined as
(11) $\tau_{i}(\mathbf{x}_{a},\mathbf{x}_{b})=\sigma_{i}\,\int_{\max(a,\,t_{i,0})}^{\min(b,\,t_{i,1})}K_{i}(\mathbf{x}_{t^{\prime}})\,\text{d}t^{\prime}.$
By choosing the appropriate primitive kernels, we show in Section 5.1 that this integral can be computed in closed form. A more detailed derivation can be found in the Supplemental material.
4.2. Solving Equation 7 for distance sampling
While Equation (7) is just a summation over segments that can be computed analytically (we show examples later in Section LABEL:sec:applications), in the most general case $L_{o,i}(\mathbf{x}_{t},\omega)$ hides the inscattering integral, which requires numerical evaluation using Monte Carlo sampling. Computing a Monte Carlo estimate would quickly become impractical; thus we need to sample a single segment per ray, which requires essentially sampling the distance $t$ along the ray, ideally with probability distribution function (PDF) $p(t)=\mu_{t}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$.
While our primitive-based media allows the use of delta tracking (or any other Monte Carlo-based) estimators for distance sampling, we found that we can leverage the recursive formulation of transmittance (9) and pose the problem as an iterative search problem, where we uniformly sample the transmittance with a random variable $\xi\in(0,1)$, and search for the distance $t$ so that $\xi(t)=\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$. This approach is very similar to regular tracking (stutton99regular), which we extend to handle overlapping kernels efficiently. In particular, we first search for a segment $\mathcal{S}_{k}$ so that $(1-\xi)\in[\textrm{T}_{k-1},\textrm{T}_{k}]$. Then, we invert the transmittance inside the segment, solving for $t\in[\widehat{t}_{k,0},\widehat{t}_{k,1}]$ the following equation:
(12) $\log(1-\xi)=-\sum_{i\in S_{k}}\tau_{i}\left(\mathbf{x}_{\widehat{t}_{k,0}},\mathbf{x}_{t}\right).$
Depending on the type of kernel $K_{i}(\mathbf{x})$ being used, and the amount of overlaps in segment $\mathcal{S}_{k}$ this equation may have a closed-form analytical solution. However, for cases where Equation (12) does not have a simple analytical solution (e.g. many overlapping kernels at the same segment), we rely on numerical root-finding. In particular, we use two different methods depending on the kernel: the Newton-Raphson and the bisection solvers. The former is our primary choice, since it is efficient and converges with good precision with a few iterations (1-3), given that our problem is well-conditioned. However, for certain long-tailed kernels (e.g., Gaussian) it occasionally suffers from numerical instability. In these cases, we rely in the slightly slower but more stable bisection method. Details for both methods can be found in the Supplemental.
Biased uniform distance sampling inside segments
To maximize speed, in complex assets fitted with a high number of small kernels we can avoid inversion (12) and simply sample inside each segment uniformly. While this means we sample with pdf $p(t)\neq\mu_{t}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$, in our implementation we assume that the probability distribution is close enough to it, allowing us to cancel out transmittance when computing the path throughput. This introduces a small amount of bias in exchange for a significant performance boost, which we analyze in Section LABEL:sec:forward_rendering.
5. Kernels
In the previous section, we have derived a general framework for radiative transport in primitive-based media, defined using arbitrary kernel functions. Any kernel could be used within our framework as long as 1) they have limited support or their decay is such that they can be bounded or clipped efficiently and 2) closed-form solutions to their line integrals exist or can be numerically computed efficiently.
Here we implement our framework using two different kernels: the Gaussian and the Epanechnikov kernel. These have been successfully used in the signal processing, density estimation, rendering, and inverse graphics literature (liu2021epan; jakob2011progressive; kerbl20233Dgaussians). Figure 3 shows their shape; we compare their performance in different applications in Section LABEL:sec:kernelcomp.
Figure 3. Visual comparison between the 1D (insets) and 2D Gaussian and Epanechnikov kernels. We plot $\pm{}3\sigma$ support to visualize the main difference between them: while the Gaussian kernel has infinite support, the Epanechnikov is finite, and decays sharply to 0. We further analyze differences between these kernels in Section LABEL:sec:kernelcomp.
[…]
[¡short description¿]¡long description¿
5.1. Gaussian Kernel
The normalized Gaussian kernel $g_{i}(\mathbf{x})$ for primitive $\mathcal{P}_{i}$ is defined as
(13) $g_{i}(\mathbf{x})=\frac{1}{(2\pi)^{\frac{3}{2}}(|\Sigma_{i}|)^{\frac{1}{2}}}e^{-\frac{1}{2}(\mathbf{x}-\bm{c}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\bm{c}_{i})},$
where $\Sigma_{i}$ is the $3\times 3$ covariance matrix and $\bm{c}_{i}$ its mean. As opposed to other convex kernels, Gaussians have unlimited support, which would potentially require evaluating all primitives in the scene for every ray. In practice, similar to Kerlb et al. (kerbl20233Dgaussians), we bound the kernel’s support to three times the standard deviation, covering 99.73% of the Gaussian’s density.
Transmittance Evaluation
We compute the closed-form transmittance of a Gaussian primitive $\mathcal{P}_{i}$ from point $\mathbf{x}_{t_{0}}$ to point $\mathbf{x}_{t_{1}}$ along the ray $\mathrm{r}(t)=\mathbf{x}_{0}+\omega\,t$, by integrating the optical depth $\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})$. These parameters are them defined in the Gaussian’s local coordinate system, also known as whitened space (which we obtain by linearly transforming the ray into the canonical space of the Gaussian ellipsoidal shell, where the ellipsoid is a sphere of radius 1). Due to the symmetry of the Gaussian kernel around its mean, we can simplify the definite integral by shifting the origin of the ray to ensure a symmetric integration domain, integrating optical depth $\tau_{i}(\mathbf{x}_{-t_{s}},\mathbf{x}_{t_{s}})$ with $t_{s}=\frac{t_{0}+t_{1}}{2}$. This reduces the number of error functions required from 2 to 1. Plugging in Equation (13) into Equation (11) we get
(14) $\tau_{i}(\mathbf{x}_{-t_{s}},\mathbf{x}_{t_{s}})$ $=\sigma_{i}\int_{-t_{s}}^{t_{s}}g_{i}(\mathbf{x}_{t})\,\text{d}{t}=Ge^{-\frac{1}{2}\left(a-b^{2}\right)}\left[\operatorname{erf}\left(t_{s}\sqrt{\frac{1}{2}}\right)\right]$
with
(15) $a=\mathbf{x}_{0}\times\mathbf{x}_{0},\quad b=\mathbf{x}_{0}\times\omega,\quad G=\frac{\sigma_{i}}{2\pi\sqrt{v^{T}\Sigma_{i}^{-1}v|\Sigma_{i}|}}$
with $\operatorname{erf}(\cdot)$ being the error function. Assuming exponential media, we can compute the transmittance as $T(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})=\exp(-\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}}))$.
Distance Sampling
For sampling the free-flight distance $t$ in a single Gaussian primitive $\mathcal{P}_{i}$ with PDF $p(t)=\alpha_{i}G_{i}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$, we need to invert its optical depth (14), which by setting $t_{0}=0$ and using a random value $\xi\in(0,1)$, and assuming an accumulated transmittance up to $t_{0}$ of $\beta$ has closed form following
(16) $t(\xi)=\sqrt{2}\operatorname{erf}^{-1}\left(\operatorname{erf}\left(\frac{b}{\sqrt{2}}\right)-\frac{2\operatorname{log}(\xi/\beta)}{Ge^{-\frac{1}{2}(a-b^{2})}}\right)-b$
For segments where more than one 3D Gaussian is contributing, we revert to using the bisection solver presented in Section 4. For single kernels, we normally use Equation 16. In practice, the piece-wise approximation of the $\operatorname{erf}^{-1}$ function commonly implemented in GPU math libraries (cuda) produces large errors for certain input values, and for these we resort to the bisection solver as well, at the cost of some performance.
5.2. Epanechnikov Kernel
In our primitive-based rendering framework, Epanechnikov kernels have a significant advantage compared to Gaussian kernels due to their limited support. This allows more compact primitive shells, which accelerate rendering by reducing the number of overlaps. The 3D Epanechnikov kernel $\mathcal{E}_{i}(x)$ for primitive $\mathcal{P}_{i}$ is defined as
[…]
where, analogously to the Gaussian kernel, $\Sigma_{i}$ is the $3\times 3$ covariance matrix, $\bm{c}_{i}$ the mean, and $d(\mathbf{x})=(\mathbf{x}-\bm{c}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\bm{c}_{i})$.
Transmittance Evaluation
We derive the closed-form transmittance of an Epanechnikov-based primitive between points $\mathbf{x}_{t_{0}}$ and $\mathbf{x}_{t_{1}}$ by integrating the optical depth between them as
(19) $\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})$ $=\sigma_{i}\int_{t_{0}}^{t_{1}}\mathcal{E}_{i}(\mathbf{x}_{t})\text{d}{t}$
$=\sigma_{i}{}\mathcal{K}_{\text{norm}}(\mathcal{K}_{3}{}(t_{0}^{3}-t_{1}^{3})+\mathcal{K}_{2}{}(t_{0}^{2}-t_{1}^{2})+\mathcal{K}_{1}{}(t_{0}-t_{1})),$
where $\mathcal{K}_{1}$, $\mathcal{K}_{2}$, $\mathcal{K}_{3}$ and $\mathcal{K}_{\text{norm}}$ are constants that depend on the ray position and direction, and the eigenvalues of the kernel’s covariance matrix $\Sigma_{i}$. Explicit expressions for these constants can be found in Appendix LABEL:app:kernel_transmittance. Analogously to any other kernel, and assuming exponential media, we can finally compute the transmittance as $T(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})=\exp(-\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}}))$.
Distance Sampling
While there is an analytic solution for inverting Equation (12) for one Epanechnikov kernel, it is not practical, given its complexity. Instead, we directly use the Newton-Raphson solver for segments with both one or multiple overlapping kernels.
6. Implementation Details
We implement our method in Mitsuba 3 (jakob2022mitsuba3), which we extended to support volumes of kernel primitives with different shell geometries, where we use ray tracing for querying primitives along the ray. We implement two different integrators, depending on the target application: 1) a volumetric-primitives path tracer (VPPT) supporting scattering media, and 2) a simplified volumetric-primitives radiance field (VPRF) integrator that computes the radiance field along a primary ray.
In both cases, we compute transmittance and emission by following the segment-based formulation described in Section 4.2, by iterating over the segments in an ordered fashion: This fits very well with our ray-tracing-based querying of the primitives, since we collect new segments by simply casting new rays from the previous intersection. Listing LABEL:code:primitive-tracing shows a skeleton pseudocode for our two integrators.
6.1. Integrators
Volumetric-primitives path tracer (VPPT)
This integrator is mostly an off-the-shelf volumetric path tracer supporting next-event estimation, with the key modifications of the transmittance evaluation and sampling, and the medium interaction routines. It solves Equation (7) using Monte Carlo estimation over multiple paths where, assuming non-emissive media, the throughput of each sample path is recursively computed as
(20) $L(\mathbf{x}_{0})=\beta(\mathbf{x}_{0},\mathbf{x}_{i})\mu_{t}(\mathbf{x}_{i})\alpha(\mathbf{x}_{i})S(\mathbf{x}_{i},\omega_{i}),$
with $\omega_{i}=\frac{\mathbf{x}_{i}-\mathbf{x}_{i-1}}{|\mathbf{x}_{i}-\mathbf{x}_{i-1}|}$, and $\beta(\mathbf{x}_{0}\rightarrow\mathbf{x}_{i})$ the throughput of the eye subpath until $\mathbf{x}_{i}$ defined as
(21) $\beta(\mathbf{x}_{0},\mathbf{x}_{i})=\frac{\beta(\mathbf{x}_{0},\mathbf{x}_{i-1})\textrm{T}(\mathbf{x}_{i-1},\mathbf{x}_{i})\mu_{t}(\mathbf{x}_{i-1})\alpha(\mathbf{x}_{i})f_{\text{p}}(\mathbf{x}_{i-1},\omega_{i-1}\rightarrow\omega_{i})}{p(\mathbf{x}_{i})},$
where $p(\mathbf{x}_{i})$ is the probability of generating $\mathbf{x}_{i}$ from the previous scattering vertex $\mathbf{x}_{i-1}$. In our implementation, we obtain $\mathbf{x}_{i}$ using the standard practice of first sampling the phase function at $\mathbf{x}_{i-1}$, and then sampling distance with probability proportional to $\textrm{T}(\mathbf{x}_{i-1},\mathbf{x}_{i})$.
Volumetric-primitives radiance field (VPRF)
This implements in our volumetric primitives framework the radiance field image formation model, similar to e.g. (mildenhall2020; kerbl20233Dgaussians). It significantly simplifies the image formation model from VPPT, by only considering the primary rays from the camera, accumulating radiance along the ray by querying a spherical harmonics-based emission in the primitives. In addition, we disable overlapping logic in order to favour speed. In practice, this means that we approximate Equation (7) using an ad-hoc approximate of the emission at each segment following
[…]
6.2. Ray tracing volumetric primitives
As described above, we limit the support of our kernels by bounding their bandwidth with an ellipsoid shell. We query the primitives along a ray leveraging ray-tracing, which we accelerate through the construction of a BVH using Optix (optix) for efficient ray traversal. We implemented a ray-ellipsoid intersection as a custom intersection routine in Optix. However, we found that this produces suboptimal BVHs given the potentially large support of the primitives. Moreover, a custom intersection kernel is significantly less performant than a hardware-accelerated ray-triangle intersection test. Thus, instead of using the ray-ellipsoid intersection, we can triangulate the ellipsoids, and ray-trace on triangles. We analyzed a variety of triangulation approaches (see Figure LABEL:fig:shell_ablation), and found a significant boost on performance despite the higher number of primitives in the scene, finding highly-tesselated icospheres to perform best in terms of execution times. One could consider using mesh instancing to avoid duplicating the shell’s geometry. However, this approach reintroduces the problem of using axis-aligned bounding boxes for the individual instances in the instances BVH.
Stack array allocation
A limitation of our primitive tracing algorithms is their need to stack information while iterating over the primitives along the ray. This is particularly necessary for calculating the list of primitive exit points during segment iteration, as well as maintaining a record of the active primitives for the current segment. Unfortunately, most JIT frameworks, such as Dr.Jit or PyTorch, do not offer an API with a low enough level to perform pointer arithmetic. For efficiency, it is crucial to execute this performance-critical algorithm in registers, which necessitates the ability to allocate, access, and write into arrays of variables on the stack rather than on the heap. Fortunately, the NVidia PTX assembly language offers local-state space private memory for each thread to store its own data, which can be utilized for this purpose. We modified the Dr.Jit (jakob2020drjit) core to include routines in the JIT compiler that generate such intrinsics.